When the length of a rectangular piece of sheet metal is 2 feet more than the width its area is 35 square feet. What are the length and width of this piece of sheet metal?
Answer by mananth(16946)

Area = 35 sq. ft
Width = x ft
Length =(x+2)
Area =x *1(x+2)
35 =x^2+2x
x^2+2x-35=0
(x+7)(x-5)=0
x=5 the positive value
width = 5 and length =7 feet